OpenStudy (anonymous):

please help solve the system using substitution, explain your steps for full credit. actually want to learn the steps not just the answer. 3x-4y=10 2y=2x-6 medal given

OpenStudy (mrdoe):

pick whichever looks simpler to you and solve for one variable in terms of the other. then substitute that into the other equation to explicitly solve for one variable.

OpenStudy (phi):

using substitution means "solve" for one of the variables as a first step. Example: 2y=2x-6 (is one of your equations) what happens if we divide both sides by 2 (that means divide *all terms* by 2) 2y/2 = 2x/2 - 6/2 can you simplify each term ?

OpenStudy (phi):

can you simplify \[ \frac{2y}{2} = \frac{2x}{2} - \frac{6}{2} \] or is it a mystery ?

OpenStudy (phi):

can you simplify 6/2 ?

OpenStudy (anonymous):

thank you. I now understand,

OpenStudy (phi):

when you get the answer, please post it.

OpenStudy (anonymous):

I sure will, have to go to work, but will work on it tomorrow.p.s to finish the problem I just plug the answer back into the other problem?

OpenStudy (phi):

"just plug" yes, it is easy if you know how to do this... but hard if you don't. And (because it was not obvious to you how to solve for y in the 2nd equation) I am guessing it is a difficult problem for you. But post what you get and we will help you.

OpenStudy (anonymous):

thanks